a) Check whether the matrices A and B are diagonalisable. Diagonalise those
matrices which are diagonalisable.
A = 1 0 0 B = 2 0 0
1 5 -3 -2 2 -1
2 8 -5 -1 0 1
b) Find inverse of the matrix B in part a) of the question by using
Cayley-Hamiltion theorem.

c) Find the inverse of the matrix A in part a) of the question by fnding its
adjoint.
1
Expert's answer
2020-07-13T19:50:08-0400

a) Find eigenvalues from the characteristic polynomial of A:

"\begin{vmatrix} 1-\lambda & 0 & 0 \\ 1 & 5-\lambda & -3 \\ 2 & 8 & -5-\lambda \end{vmatrix}" "=\u2212\u03bb^3+\u03bb^2+\u03bb\u22121=\u2212(\u03bb\u22121)\u00d7(\u03bb^2\u22121)=\u2212(\u03bb\u22121)\u00d7(\u03bb\u22121)\u00d7(\u03bb+1)"

Hence, Eigenvalues are 1, 1, -1.

Corresponding to eigenvalue 1, G.M. = "3 - \rho(A-I)"

= "3 - \rho \begin{bmatrix} 0 & 0 & 0 \\ 1 & 4 & -3 \\ 2 & 8 & -6 \end{bmatrix} = 3-1 = 2" .

Corresponding to eigenvalue -1, G.M. = 1.

Hence, A matrix is diagonlizable.


Characteristic polynomial of B is

"\begin{vmatrix} 2-\lambda & 0 & 0 \\ -2 & 2-\lambda & -1 \\ -1 & 0 & 1-\lambda \end{vmatrix}=0"

"\implies -\lambda^3+5 \lambda^2-8\lambda+4 = 0 \\
=-(\lambda-1)(\lambda^2-4\lambda+4)=-(\lambda-1)(\lambda-2)(\lambda-2) = 0"

Hence, eigenvalues of B are 1,2,2.

Corresponding to eigenvalue 2, G.M. = "3-\rho(A-2I) = 3 - \rho \begin{bmatrix} 0 & 0 & 0 \\ -2 & 0 & -1 \\ -1 & 0 & -1 \end{bmatrix}"

= 3 - 1 = 2.

Hence, B matrix is not diagonalizable.


b) According to Cayley-Hamilton theorem, "Ch(B)=0" .

"-B^3+5 B^2-8B+4I = 0 \\
\implies 4I = B^3 - 5B^2+8B \\
\implies B^{-1} = \frac{1}{4} [B^2-5B+8I]"

So, "4 B^{-1} = \begin{bmatrix} 2 & 0 & 0 \\ -2 & 2 & -1 \\ -1 & 0 & 1 \end{bmatrix} \begin{bmatrix} 2 & 0 & 0 \\ -2 & 2 & -1 \\ -1 & 0 & 1 \end{bmatrix} - 5 \begin{bmatrix} 2 & 0 & 0 \\ -2 & 2 & -1 \\ -1 & 0 & 1 \end{bmatrix} + 8 \begin{bmatrix} 1 & 0 & 0 \\ 0 & 1 & 0 \\ 0 & 0 & 1 \end{bmatrix}"

"\\implies B^{-1} = \\begin{bmatrix} \\frac{1}{2} & 0 & 0 \\\\ \\frac{1}{4}\t& \\frac{1}{2}\t& \\frac{-1}{2} \\\\ \n\\frac{1}{2}\t & 0 & \t1 \\end{bmatrix}" .


c) "|A| = -25+24 = -1" ,

Cofactors of A are "C_{11} = -2, C_{12} =-1, C_{13} = -2, C_{21} = 0, C_{22} = -5, \\
C_{23} =-8, C_{31} = 0, C_{32} = 3, C_{33} = 5"

"adj(A) = \\begin{bmatrix} -1 & -1 & -2 \\\\ 0\t& -5\t& -8 \\\\ \n0 & 3 & \t5 \\end{bmatrix}^T = \\begin{bmatrix} -1 & 0 & 0 \\\\ -1\t& -5\t& 3 \\\\ \n-2\t & -8 & \t5 \\end{bmatrix}"

"A^{-1} =\\frac{1}{|A|}adj(A) = \\begin{bmatrix} 1 & 0 & 0 \\\\ 1\t& 5\t& - 3 \\\\ 2\t & 8 & \t-5 \\end{bmatrix}" .